OpenStudy (anonymous):

Find the Laplace transform of the following function. f (t) = t^3 H( t − 2) + e^4t H (t − 3)

OpenStudy (unklerhaukus):

\[f (t) = t^3 H( t − 2) + e^{4t} H (t − 3)\]\[\mathcal L \{f(t)\}=\]

OpenStudy (foolaroundmath):

\[\Large L(f(t-\alpha).H(t-\alpha)) = \int_{0}^{\infty}e^{-st}.f(t-\alpha).H(t-\alpha)dt\]\[\Large = \int_{\alpha}^{\infty}e^{-st}f(t-\alpha)dt\] Taking \(z = t-\alpha \) \[\Large = \int_{0}^{\infty}e^{-s(z+\alpha)}f(z)dz = e^{-s\alpha}\int_{0}^{\infty}e^{-sz}f(z)dz \]\[\Large =e^{-s\alpha}L(f(t)) \]So, what you want to do is to express the given function as a sum of f(t-a).H(t-a) and then apply this relation

OpenStudy (turingtest):

H is the heavyside function I take it?

OpenStudy (unklerhaukus):

i think so

OpenStudy (turingtest):

\[\mathcal L\{u_c(t)g(t)\}=e^{-cs}\mathcal L\{g(t+c)\}\]may work

OpenStudy (turingtest):

\(u_c\) is how I write the heaviside function

OpenStudy (unklerhaukus):

my book uses \[h(t − a) \]

OpenStudy (turingtest):

oh for the heaviside, nvmd

OpenStudy (turingtest):

same thing\[h(t-a)=h_a(t)\]

OpenStudy (unklerhaukus):

yerh for the unit heavyside step function

OpenStudy (foolaroundmath):

these set of lectures (specially the second one might help you with this problem) http://home.iitk.ac.in/~sghorai/TEACHING/MTH203/ode17.pdf http://home.iitk.ac.in/~sghorai/TEACHING/MTH203/ode18.pdf http://home.iitk.ac.in/~sghorai/TEACHING/MTH203/ode19.pdf

OpenStudy (anonymous):